# Shadow of Quantum Improved Regular Kerr Black Hole and parameter constrains with EHT observations

### v1

@@ VI Constraints from EHT Observations
$R(\varphi)=\sqrt{(\alpha-\alpha_{C})^{2}+(\beta-\beta_{C})^{2}},\quad\tan(\varphi)=\frac{\beta-\beta_{C}}{\alpha-\alpha_{C}},$ (45)
where $(\alpha_{C},\beta_{C})$ represent the shadow’s displacement from the black hole center at $(0,0)$. Given the intrinsic axisymmetry, the vertical displacement is zero ($\beta_{C}=0$), and the horizontal displacement $\alpha_{C}$ is
$\alpha_{C}=\frac{\alpha_{r}+\alpha_{l}}{2},$ (46)
where $\alpha_{r}$ and $\alpha_{l}$ denote the abscissae where the shadow intersects the $\alpha$-axis.
@@
In the following, we assume that M87* and Sgr A* are QIRK black holes and aim to constrain the model parameters based on astronomical data. By analyzing these observables, we will exclude parameter ranges incompatible with the data, using density maps based on Fig. 3. While uncertainties in the EHT observations of M87* and Sgr A* introduce some degree of imprecision, we can still estimate the approximate effect of the parameters introduced by the QIRK black holes in the observations.
In Sec. V, we analyze the black hole shadow assuming an inclination angle of $\theta_{0}=90^{\circ}$, which maximizes deviation from circularity and enhances the effects of the black hole’s properties. However, various inclination angles can be considered for observational constraints. For instance, EHT estimates the inclination angle of M87* to be approximately $163^{\circ}$, based on the relativistic jet orientation CraigWalker:2018vam . Since the shadow is symmetric about the $\alpha$-axis, $\theta_{0}=163^{\circ}$ is equivalent to $\theta_{0}=17^{\circ}$ in our analysis. For Sgr A*, while the inclination angle is not definitively known, values below $52^{\circ}$ are generally favored. Inclinations of $5^{\circ}$ Kuang:2022ojj , $50^{\circ}$ walia_testing_2022 , and $90^{\circ}$ Ghosh:2022kit have been employed in previous analyses. The circularity deviation $\Delta C$ increases with the inclination angle $\theta_{0}$, so we can obtain an upper bound on the parameters by choosing an inclination of $90^{\circ}$. However, the shadow area decreases as the inclination angle increases, and the angular diameter $\theta_{d}$ and Schwarzschild deviation $\delta$ are proportional to the area. Therefore, smaller inclination angles can provide tighter constraints when considering $\theta_{d}$ and $\delta$. After taking these factors into account, we select $\theta_{0}=17^{\circ}$ and $90^{\circ}$ for M87*, and $50^{\circ}$ for Sgr A*.
VI.1 Constrains from M87*
@@
Figures 9, 10, and 11 display the density plots of $\Delta C$, $\theta_{d}$, and $\delta$ under varying inclinations and parameters. As shown in Fig. 9, $\Delta C$ remains below 0.10 for $\theta_{0}=17^{\circ}$, providing no substantial constraints. However, for $\theta_{0}=90^{\circ}$, some parameters with $\Delta C>0.10$ are excluded. From Fig. 10, we can observe that the area of the shadow $A$ increases as the inclination angle $\theta_{0}$ decreases. Therefore, $\theta_{0}=17^{\circ}$ provides stronger constraints if we exclude $\theta_{d}<39\mu\text{as}$. For $\theta_{d}$, the exclusion criteria are $\theta_{d}>45\mu\text{as}$ or $\theta_{d}<39\mu\text{as}$ within the $1\sigma$ interval. Given the modest deviation of the QIRK black hole from the Kerr black hole, we do not extend the analysis to the $2\sigma$ interval. Finally, Fig. 11 shows that all values of $\delta$ lie within the $1\sigma$ interval, meaning no significant exclusion can be made based on $\delta$ alone.
@@ VI.2 Constrains from Sgr A*
For Sgr A*, only the angular diameter $\theta_{d}$ and deviation $\delta$ are considered, as the EHT did not provide a bound on the circularity deviation $\Delta C$. As with M87*, Fig. 12 allows us to exclude parameters where $\theta_{d}>50\mu\text{as}$. Additionally, Fig. 13 plots $\delta=-0.04$ (Keck), but this does not yield significant constraints.
@@
Figure 14: The observational bounds for M87* and Sgr A* with different $p$. The black and green curves represent the extremal case $(\widetilde{\omega}_{c})$, separated by $a_{\text{cri}}$ (the critical value for the shadow in the extremal case). The orange curve corresponds to $\widetilde{\omega}_{m}$, with the intersection of the two curves indicating $a_{\text{max}}$, the maximum value of $a$ for which both the black hole and the absence of CTCs are satisfied, as discussed in Sec. II. In the first panels, parameters to the right are excluded by the blue curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=17^{\circ}$, the red dashed curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=17^{\circ}$, and the magenta dashed curve $\Delta C=0.10,\,\theta_{0}=90^{\circ}$. In the second panel, parameters to the right are excluded by the red dashed curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=90^{\circ}$ and the blue curve $\Delta C=0.10,\,\theta_{0}=90^{\circ}$. In contrast, in the lower two panels, parameters to the left are excluded by the blue curve $\theta_{d}=50\,\mu\text{as},\,\theta_{0}=50^{\circ}$. In summary, the final parameter constraints are represented by the blue curves, which provide the most restrictive limits on the parameters. Meanwhile, the gray regions represent the final allowed parameter range.
@@ VII Conclusion
In this paper, building on the work of Ref. chen2024quantumimprovedregularkerr , we conduct a more detailed study of the QIRK black hole. Due to the difficulty of analytic discussions, we employed numerical methods and, after comprehensively considering the three conditions of resolving singularities, ensuring the existence of horizons, and avoiding closed timelike curves, we determined the parameter range allowed for the QIRK black hole. We found that for a given $p$, the quantum correction $\widetilde{\omega}$ has both upper and lower bounds ($\widetilde{\omega}_{c}$ and $\widetilde{\omega}_{m}$). This also indicates that, under these three conditions, the quantum effects cannot be too large or too small. Additionally, the rotational parameter $a$ has a maximum value less than 1, which differs significantly from the Kerr black hole, indicating that the QIRK black hole does not exhibit extremely high rotation. Moreover, we also discovered that the QIRK metric can describe a compact object with closed timelike curves but no horizon. The causality-violating signals from this region could potentially be observable by external observers, making this area worthy of further investigation.
As a strong field test, the black hole shadow can reveal both universal and unique features of different black hole models. To study the shadow of the QIRK black hole, we conducted a comprehensive analysis of the photon sphere structure in a more general context and provided the analytical expressions for the photon sphere. Through the shadow-related function, we constructed an intuitive image of the photon sphere structure, establishing a specific correspondence between the photon sphere and the black hole shadow. We also specifically considered the photon sphere of the QIRK black hole in both non-extremal and extremal cases, and compared it to the Kerr black hole. From the plots of the shadow-related function for the non-extremal case, we observe that their behaviors are quite similar, indicating that their shadows differ only in size. In particular, in the extremal case, the photon sphere at the horizon exhibits distinctive properties, which are reflected in the shadow.
The introduction of the quantum correction parameter $\widetilde{\omega}$ extends the extremal black hole condition from $a=1$ in Kerr black holes to a curve $\widetilde{\omega}=\widetilde{\omega}_{c}$ in the QIRK black hole. This gives the extremal QIRK black hole a richer structure, manifested in the fact that when $a<a_{\text{cri}}$, the shadow of the extremal QIRK black hole resembles that of a non-extremal black hole, making them indistinguishable. Only when $a>a_{\text{cri}}$ does the shadow of the extremal QIRK black hole resembles that of the extremal Kerr black hole. Regardless of whether the black hole is in an extremal state, the trends in shadow changes with $p$ and $\widetilde{\omega}$ remain the same, differing only in magnitude. The overall changes remain small, indicating that the quantum effects introduced by the QIRK model do not produce significant alterations in the black hole shadow.
When considering the recent observational results of M87* and Sgr A* from the EHT to further investigate the QIRK black hole as a potential candidate for astrophysical observations, we find that, due to the relatively small impact of quantum effects on the shadow, even when considering all allowed parameters, the three observables reflecting the shadow size—$\Delta C$, $\theta_{d}$, and $\delta$— still exhibit variations within a relatively small range. To obtain stronger constraints, we consider different inclination angles. For the case of $p=4$, the permissible parameters lie within the region to the left of the blue curve in Fig. 14a, and to the right of the blue curve in Fig. 14c. Both regions are shaded in grey in the figure. The final result is the intersection of these two grey regions, forming a band-like area. A similar region is formed for $p=6$, following the same pattern. Thus, based on the current results, the QIRK black hole remains a viable candidate for observational testing.
In this paper, we only discussed the cases of $p=4$ and $p=6$. This choice was made for two reasons. First, a larger $p$ increases the degree of the polynomial, making the analysis more complicated. Second, as $p$ increases, the shadow of the QIRK black hole gradually converges to that of the Kerr black hole, making them indistinguishable. Moreover, current observational data would not be sufficient to constrain the parameters in such cases. Therefore, we focused on the two smallest values of $p$ to highlight the distinctive features of this black hole model and to obtain stronger constraints.
Whether the QIRK black hole can serve as a viable candidate for a real astrophysical black hole will ultimately depend on more precise and abundant observational data. Future observations will be crucial in either verifying or ruling out the QIRK black hole as a realistic black hole model.

### v2

@@ V Constraints from EHT Observations
$R(\varphi)=\sqrt{(x-x_{C})^{2}+(y-y_{C})^{2}},\quad\tan(\varphi)=\frac{y-y_{C}}{x-x_{C}},$ (42)
where $(x_{C},y_{C})$ represent the shadow’s displacement from the black hole center at $(0,0)$. Given the intrinsic axisymmetry, the vertical displacement is zero ($y_{C}=0$), and the horizontal displacement $x_{C}$ is
$x_{C}=\frac{x_{r}+x_{l}}{2},$ (43)
where $x_{r}$ and $x_{l}$ denote the abscissae where the shadow intersects the $x$-axis.
@@
In the following, we aim to constrain the model parameters based on astronomical data. By analyzing these observables, we will exclude parameter ranges incompatible with the data, using density maps based on Fig. 3. While uncertainties in the EHT observations of M87* and Sgr A* introduce some degree of imprecision, we can still estimate the approximate effect of the parameters introduced by the QIRK black holes in the observations.
The viewing inclination generally differs from source to source. For instance, EHT estimates the inclination angle of M87* to be approximately $163^{\circ}$, based on the relativistic jet orientation CraigWalker:2018vam . Since the shadow is symmetric about the $x$-axis, $\theta_{0}=163^{\circ}$ is equivalent to $\theta_{0}=17^{\circ}$ in our analysis. For Sgr A*, while the inclination angle is not definitively known, values below $52^{\circ}$ are generally favored. Inclinations of $5^{\circ}$ Kuang:2022ojj , $50^{\circ}$ walia_testing_2022 , and $90^{\circ}$ Ghosh:2022kit have been employed in previous analyses. The circularity deviation $\Delta C$ increases with the inclination angle $\theta_{0}$, so we can obtain an upper bound on the parameters by choosing an inclination of $90^{\circ}$. However, the shadow area decreases as the inclination angle increases, and the angular diameter $\theta_{d}$ and Schwarzschild deviation $\delta$ are proportional to the area. Therefore, smaller inclination angles can provide tighter constraints when considering $\theta_{d}$ and $\delta$. After taking these factors into account, we select $\theta_{0}=17^{\circ}$ and $90^{\circ}$ for M87*, and $50^{\circ}$ for Sgr A*.
V.1 Constraints from M87*
@@
Figs. 10, 11, and 12 display the density plots of $\Delta C$, $\theta_{d}$, and $\delta$ under varying inclinations and parameters. As shown in Fig. 10, $\Delta C$ remains below 0.10 for $\theta_{0}=17^{\circ}$, providing no substantial constraints. However, for $\theta_{0}=90^{\circ}$, some parameters with $\Delta C>0.10$ are excluded. From Fig. 11, we can observe that the area of the shadow $A$ increases as the inclination angle $\theta_{0}$ decreases. Therefore, $\theta_{0}=17^{\circ}$ provides stronger constraints if we exclude $\theta_{d}<39\mu\text{as}$. For $\theta_{d}$, the exclusion criteria are $\theta_{d}>45\mu\text{as}$ or $\theta_{d}<39\mu\text{as}$ within the $1\sigma$ interval. Given the modest deviation of the QIRK black hole from the Kerr black hole, we do not extend the analysis to the $2\sigma$ interval. Finally, Fig. 12 shows that all values of $\delta$ lie within the $1\sigma$ interval, meaning no significant exclusion can be made based on $\delta$ alone.
@@ V.2 Constraints from Sgr A*
For Sgr A*, only the angular diameter $\theta_{d}$ and deviation $\delta$ are considered, as the EHT did not provide a bound on the circularity deviation $\Delta C$. As with M87*, Fig. 13 allows us to exclude parameters where $\theta_{d}>50\mu\text{as}$. Additionally, Fig. 14 plots $\delta=-0.04$ (Keck), but this does not yield meaningful constraints.
@@
Figure 15: The observational bounds for M87* and Sgr A* with different $p$. The black and orange curves correspond to $\widetilde{\omega}_{c}$ and $\widetilde{\omega}_{m}$, respectively, as in Fig. 3. Their intersection defines $a_{\text{max}}$, the maximum value of $a$ for which both the existence of a horizon and the absence of CTCs are ensured, as discussed in Sec. II. In the first panels, parameters to the right are excluded by the blue curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=17^{\circ}$, the red dashed curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=17^{\circ}$, and the magenta dashed curve $\Delta C=0.10,\,\theta_{0}=90^{\circ}$. In the second panel, parameters to the right are excluded by the red dashed curve $\theta_{d}=39\,\mu\text{as},\,\theta_{0}=90^{\circ}$ and the blue curve $\Delta C=0.10,\,\theta_{0}=90^{\circ}$. In contrast, in the lower two panels, parameters to the left are excluded by the blue curve $\theta_{d}=50\,\mu\text{as},\,\theta_{0}=50^{\circ}$. In summary, the blue curves give the tightest combined constraints; grey regions remain allowed.
@@
VI Image of Closed Timelike Curves
Closed timelike curves (CTCs) are pathological features in certain solutions of general relativity, most notably in the Kerr spacetime. The existence of CTCs is generally confined within the Cauchy horizon ($r<r_{-}$), where the metric component $g_{\phi\phi}$ can become negative, theoretically violating causality. Interestingly, the parameter space of the QIRK model reveals cases in which the horizon existence condition $\widetilde{\omega}\leq\widetilde{\omega}_{c}$ is not met, and simultaneously the condition to avoid CTCs $\widetilde{\omega}>\widetilde{\omega}_{m}$ is also not satisfied. This gives rise to a theoretical possibility of ”naked CTCs” region containing closed timelike curves that are not shielded by a horizon (corresponding to the green-shaded area in Fig. 3). Although the physical realization of such spacetimes is highly speculative and likely disfavored, they are permitted within the QIRK parameter space. This prompts us to investigate the optical appearance and light propagation characteristics in this region. Such an analysis helps us understand how quantum corrections influence light propagation within this region, thereby further revealing the impact of quantum effects on causal structure.
In the following discussion, we fix the black hole spin parameter to $a=1.05$ and investigate the effects of quantum corrections by varying the quantum correction parameter $\widetilde{\omega}$. When $\widetilde{\omega}=0$, this corresponds to the Kerr naked singularity. For $0<\widetilde{\omega}\leq\widetilde{\omega}_{m}$, the solution represents a compact object.
Figure 16: The cross section at $\phi=0$ of the region where closed timelike curves exist, with the boundary defined by $g_{\phi\phi}=0$. Here we fix the spin parameter at $a=1.05$ and set $p=4$. Inside this boundary, $g_{\phi\phi}<0$, indicating the presence of closed timelike curves along the $\phi$ coordinate lines. As the quantum correction parameter $\widetilde{\omega}$ increases, this region steadily shrinks and retreats away from $r=0$.
As shown in Fig. 16, the $\phi=0$ cross section displays the boundary defined by $g_{\phi\phi}=0$ in both the Kerr spacetime and QIRK spacetimes with varying quantum correction strength $\widetilde{\omega}$. Since we only consider closed timelike curves along the $\phi$ coordinate lines, the region inside this boundary corresponds to where such curves exist. In the Kerr case (red curve), the region in which $g_{\phi\phi}$ becomes negative typically reaches the ring singularity. However, when quantum corrections are introduced, this picture changes significantly. Even a small quantum correction (e.g., $\widetilde{\omega}=0.001$, blue curve) causes the CTC region’s boundary to shift slightly outward, away from $r=0$, compared to Kerr. As $\widetilde{\omega}$ increases, the CTC region moves outward and shrinks overall, and this trend becomes more pronounced. When $\widetilde{\omega}$ is large enough (orange curve), the CTC region is substantially reduced, and its boundary is clearly distanced from $r=0$. Thus, we see that the inclusion of quantum corrections not only regularizes the ring singularity, but also significantly modifies the geometry in the immediate neighbourhood of $r=0$ (the Kerr ring singularity).
Figure 17: Images of CTCs located at $r=-0.5$ and $\theta=80^{\circ}$ are shown for various spacetime parameters. The black dashed curve marks the turning point at $r=0$. Light emitted from $r<0$ can only reach the region within this boundary. Solid curves represent the direct image and the lensed ring associated with CTCs. Dashed curves correspond to the related photon ring features. Different colors indicate different spacetime models: red corresponds to the Kerr spacetime, blue to the QIRK spacetime with $\widetilde{\omega}=0.001$, magenta to $\widetilde{\omega}=0.01$, and orange to $\widetilde{\omega}=0.1$.
Figure 18: The plot shows the radial function $R(r)$ as a function of $r$, with points A, B, and C corresponding to those marked in Fig. 17. It can be seen that for light rays reaching points outside the black dashed curve in Fig. 17, the turning points lie in the region $r>0$, which prevents the image from extending beyond the dashed boundary. The black dashed line and red point mark the location of the closed timelike curves at $r=-0.5$.
Having identified the CTC region under different quantum corrections, we now turn to examining how these corrections influence the observed images of light emitted from the region. To this end, we consider a luminous curve located at $r=-0.5$, $\theta=80^{\circ}$, and $t=0$, which lies along a closed timelike curve inside the previously discussed CTC region. Fig. 17 presents the image of this source as seen by a distant observer at $r=100$, $\theta_{o}=85^{\circ}$, and $\phi=0$. The black dashed curve consists of points with $\eta=0$ for which $R(r)=R^{\prime}(r)=0$. It is not hard to find that $R^{\prime}(0)$ always vanishing for any $\tilde{\omega}$, $M$, $a$, $\eta$, and $\xi$. Besides, the condition $R(0)=0$ is independent of $\tilde{\omega}$. Therefore, as shown in the figure, the position of the black dashed curve is independent of $\tilde{\omega}$. Any light reaching the observer’s screen that was emitted from the CTC region must lie inside this boundary. This is confirmed in Fig. 18, where a point outside the black dashed curve, such as point C, has a turning point at $r>0$, indicating that such light cannot travel from $r<0$ to the observer. Inside this dashed boundary, solid curves represent both the direct and lensed images of the source, while dashed curves mark the photon ring image that originate from the unstable spherical orbits lying at $r<0$ (present in both the Kerr naked singularity and the QIRK compact object). The corresponding impact parameters for these orbits are those plotted in Fig.4.
In the Kerr naked singularity, the direct image and lensed image (solid line) is relatively compact and confined. Even with a small quantum correction in the QIRK model, the direct image on the observer’s screen shows a modest outward expansion, and the expansion grows with $\widetilde{\omega}$. For $\widetilde{\omega}=0.1$, the direct image extends noticeably further than in the Kerr case, indicating that the quantum corrected geometry allows light from the negative-$r$ region to reach the observer over a wider angle. Meanwhile, the photon ring structure (dashed lines) also changes significantly with increasing $\widetilde{\omega}$. In the Kerr case, the photon ring associated with the $r=-0.5$ source forms a relatively tight boundary. In QIRK spacetimes, this curve shifts outward and becomes flatter as $\widetilde{\omega}$ increases. This behavior suggests that quantum corrections not only affect the propagation of directly emitted light, but also fundamentally modify the unstable photon orbits near $r=0$. Taken together, the changes in both direct and higher-order images show how the QIRK model’s regularization of the $r=0$ region alters the observable features of nearby sources.
Figure 19: Left panel: Initial azimuthal angle $\phi_{i}$ of photons emitted from the CTCs as a function of the observed horizontal coordinate $x$. Right panel: Arrival time $t_{f}$ of these photons at the observer as a function of $x$. Different colors correspond to different spacetime parameters as shown in Fig. 17.
Figure 20: Left panel: Two bundles of rays originate from a closed timelike curve at $r=-0.5$, $\theta_{s}=80^{\circ}$. The solid curves show directly emitted photons, while the red dashed curve traces the rays that form the photon ring image. Along the CTC (light-blue), the red, blue, magenta, and orange segments indicate the region visible to the observer for the Kerr case and for $\widetilde{\omega}=0.001$, $0.01$, and $0.1$, respectively–visibility here refers only to the direct and lensed images. Right panel: These rays cross the equatorial disk at $r=0$ (dark gray) and propagate outward to the observer; the black circle indicates the disk’s outer boundary.
The left panel of Fig. 19 shows the relation between the initial emission angle $\phi_{i}$ at the source and the horizontal screen coordinate $x$ for the direct and lensed images (solid curves in Fig.17). As $\widetilde{\omega}$ increases, the range of $x$ expands, consistent with the changes in Fig. 17. For any fixed $x$, the required emission angle $\phi_{i}$ shifts significantly with $\widetilde{\omega}$, indicating that quantum corrections alter the mapping from source angle to image position. This shift arises from modified null geodesics in the QIRK spacetime and reflects changes in gravitational lensing for $r<0$. Each curve also develops a small ”hook” at its right end, indicating that one emission angle $\phi$ can produce two observed $x$ values—this corresponds to the lensed image. Moreover, the observer cannot receive photons emitted at all $\phi$ angles. The range and location of observable $\phi$ angles grow with $\widetilde{\omega}$, moving toward $2\pi$.
Figure 20 illustrates two families of null rays emitted from the CTC at $r=-0.5$ and traversing the $r=0$ surface before reaching the observer. The solid curves trace the directly emitted photons, while the red dashed curves correspond to rays that form the photon ring image. In the left panel of Fig. 20, the red, blue, magenta, and orange segments mark the emission angles visible to the observer for $\widetilde{\omega}=0$, $0.001$, $0.01$, and $0.1$, respectively, matching the behavior seen in Fig. 19. The region of observable emission angles also depend on the source radius $r_{s}$, the source polar angle $\theta_{s}$, and the observer’s coordinates $r_{o}$ and $\theta_{o}$. Here we hold both the source and the observer at fixed coordinates $(r,\theta)$ while examining the effects of quantum corrections.
It is worth noting that we have omitted the photon ring image here: the photon ring probes only the underlying $r<0$ geometry and arises from unstable spherical orbits that can be launched from any azimuthal angle $\phi_{i}$, so it does not carry additional information about the closed timelike curve itself. For the same reason, our subsequent analysis of photon arrival times $t_{f}$ focuses exclusively on the direct and lensed images (the solid curves in Fig. 17).
The right panel of Fig. 19 shows the arrival time $t_{f}$ of these photons at the distant observer as a function of their observed horizontal coordinate $x$. For QIRK spacetimes with small quantum corrections, the arrival times shift only slightly compared to the Kerr case, and the shape of the curve remains largely unchanged. For larger quantum corrections, the arrival time curve changes markedly: photons arrive earlier than in the other cases. At the right end of each curve a hook-like feature appears, corresponding to the lensed image arrival times shown in the left of Fig. 19. For example, for $\widetilde{\omega}=0.1$, all points on the CTC emit rays simultaneously at t=0, and these rays are observed at different time by the observer. A bright spot firstly appears at the left endpoint of the orange curve. Then it moves to the right in the image. At the time corresponding to the other end of the orange curve, another bright spot appears and moves to left. Eventually, the two bright spots converge and vanish. The clear shift in arrival times with increasing $\widetilde{\omega}$ highlights the strong influence of quantum corrections on the underlying spacetime geometry.
@@ VII Conclusion
In this work, we focus on the basic properties of the QIRK black hole, including its parameter space, its observational signatures in thin disk imaging, parameter constraints from EHT data, and a preliminary exploration of the region containing closed timelike curves.
First, by analyzing the conditions for spacetime regularity, horizon existence, and the absence of CTCs, we identified the allowed parameter range of $(a,\widetilde{\omega})$. To ensure regularity, the exponent $p$ in the running Newton constant $G(r)$ must be an even integer greater than three. Horizon existence requires $\widetilde{\omega}\leq\widetilde{\omega}_{c}(a)$, while avoiding CTCs demands $\widetilde{\omega}>\widetilde{\omega}_{m}(a)$. Together these conditions define the physically viable region in Fig. 3, with the maximum spin satisfying $a_{\text{max}}<1$.
Next, we examine the optical appearance of a QIRK black hole illuminated by a geometrically thin, optically thick accretion disk. We find that the quantum correction $\widetilde{\omega}$ has only a small effect on the intensity and redshift distributions; its main impact is a reduction of the overall image brightness. Thus, separating QIRK from Kerr based solely on image morphology or brightness will be challenging and likely requires very sensitive measurements.
We then use shadow observables—the angular diameter $\theta_{d}$ and circularity deviation $\Delta C$— together with EHT observations of M87* and Sgr A*, to constrain $(a,\widetilde{\omega})$. These constraints form a narrow band in parameter space. For $\widetilde{\omega}=0$, they recover known limits on the spin $a$ for Kerr. As $\widetilde{\omega}$ grows, the allowed range of $a$ shrinks. Interestingly, even nearly nonrotating QIRK black holes ($a\approx 0$) can match Sgr A* data if $\widetilde{\omega}$ lies close to its critical value $\widetilde{\omega}_{c}(0)$.
Finally, we conduct a preliminary study of the optical characteristics of region associated with closed timelike curves in QIRK spacetime. Quantum corrections alter light paths in this region: the direct image becomes asymmetric, and both direct and secondary images expand outward as $\widetilde{\omega}$ increases. Photon arrival times also shift, highlighting the effect of quantum corrections on deep interior geometry.
In summary, the QIRK model offers a regular alternative to Kerr with distinctive observational features. Although some effects are subtle, variations in image brightness and EHT constraints provide key tests of this quantum gravity–inspired solution. Future high resolution observations and multiwavelength data should tighten limits on $\widetilde{\omega}$ and $a$. On the theoretical side, complex astrophysical environments (e.g., thick accretion disks and jets) and detailed analyses of thermodynamic properties, stability, and quasinormal modes Stashko:2024wuq are important aspects for future studies.